# Chiral nature of current flow in the regime of the quantized Hall effect

### v2

@@ II.1 Coherent and chiral model of $n$-terminal device
Figure 2: Simplified circuit diagram of an $n$-terminal device. The inner circle indicates coherent and chiral current flow inside the 2DES. Each contact $i=1,...,n$ has an ohmic resistance $R_{i}$ and a voltage $V_{i}$ applied in respect to the electrical ground. $V_{i}^{\prime}$ are the voltages resulting at the intersections between the ohmic contacts and the coherent inner circle. The magnetic field vector pointing into the plane of the 2DES gives rise to a right handed chirality of a free charge which translates into a left handed chirality of the device, cf. circled arrow.
The inner circle in Fig. 2 indicates the 2DES, the radial lines symbolize the $n$ terminals $i=1,2,...,n$ connected via ohmic contacts to the 2DES. External voltages $V_{i}$ can be applied to the contacts in respect to ground, while $V_{i}^{\prime}=V_{i}+I_{i}R_{i}$ are the voltages present at the intersects of the 2DES and each contact. The current $I_{i}$ flows through the $i$th contact via the ohmic resistance $R_{i}$ to ground.
@@
where $G_{ij}$ are the conductance coefficients between mutual contacts. They can be formally derived from the scattering matrix of the problem. The left hand side of Eq. (2) is a general formulation of the transmission properties of the 2DES based on the Landauer-Büttiker formalism, while the last term of Eq. (2) applies Ohm’s law to the macroscopic contacts. If the transport within the 2DES is chiral, only those $G_{i,j\neq i}$ connecting adjacent contacts can be non-zero. The direction of the chirality is linked to the direction of the magnetic field vector. In our case sketched in Fig. 2, we consider $\vec{B}$ pointing into the plane of the 2DES, such that $G_{ij}\neq 0$ for $i=j+1$. Assuming coherent transport, the scattering matrix is unitary, implying $\sum_{i}G_{ij}=\sum_{j}G_{ij}=0$, such that our coherent and chiral conductivity tensor takes the form
$G_{ij}=\nu G_{0}\left\{\begin{matrix}1\,;&i=j+1\\[4.30554pt] -1\,;&i=j\hfill\par\LTX@newpage\\[4.30554pt] 0\,;&\text{else}\hfill\par\LTX@newpage\end{matrix}\right.\,,$ (3)
@@
where $G_{0}=e^{2}/h$ and $\nu$ is the filling factor of the (spin-resolved) Landau-levels. While the assumption of coherence guaranties a unitary scattering matrix, for the validity of Eq. (3) it is actually sufficient to require the absence of momentum scattering. For the quantized plateaus of the Hall resistance $\nu$ is an integer. Inserting Eq. (3) into Eq. (2) yields a solvable linear equation system providing us with the prediction of the current into the $i$th contact
$I_{i}=\frac{V_{i}^{\prime}-V_{i}}{R_{i}}\quad\text{with}\quad V_{i}^{\prime}=\frac{1}{P_{i}}\left(\frac{S_{n}}{P_{n}-1}+S_{i}\right)\,,$ (4)
$P_{i}=\left\{\begin{matrix}\prod_{j=1}^{i}\left(1+\frac{1}{\nu G_{0}R_{j}}\right)\,;&i=1,2,...,n\\[4.30554pt] 1\,;&i=0\phantom{,2,...,n}\end{matrix}\right.$ (5)
and
$S_{i}=\sum_{j=1}^{i}P_{j-1}\frac{V_{j}}{\nu G_{0}R_{j}}\,.$ (6)
@@
A simplifying special case of some relevance is that of $V_{1}\neq 0$ but all other applied voltages $V_{i>1}=0$, which yields the current ratios
$\frac{I_{i>1}}{I_{1}}=\frac{{P_{n}}/{P_{i}}}{1-{P_{n}}/{P_{1}}}\,\frac{1}{\nu G_{0}R_{i}}\,.$ (7)
@@ II.2 Drude model of an $n$-terminal device
where the specific conductivity (defined at $B=0$) is $\sigma_{0}=en_{\text{s}}\mu$. For a laterally confined 2DES, in the steady state we expect $j_{y}=0$, which implies $E_{y}=-\mu BE_{x}$ and, consequently, $j_{x}=\sigma_{0}E_{x}=R_{\text{H}}^{-1}E_{y}$.
@@
The longitudinal component $j_{xx}=\sigma_{xx}E_{x}$ is the ohmic term, proportional to the specific conductivity and driven by the longitudinal electric field $E_{x}$. Caused by the additional Hall resistance at finite $B$, the ohmic current density is reduced by the factor $\left[1+(\mu B)^{2}\right]^{-1}$. The transversal component $j_{xy}=\sigma_{xy}E_{y}$ is driven by the perpendicular electric field and attenuated by the non-local Hall resistance. For $\mu B\gg 1$ the ohmic component $j_{xx}$ vanishes and the current density $j\to j_{xy}$ becomes independent of $\sigma_{0}$. It implies, that momentum scattering is absent. At the same time, the transversal $j_{xy}$ is bent into a current carrying contact as it follows the charged edge of the Hall bar.
@@
Figure 3: Drude model: Simplifying sketch showing the partitioning of the current contributions $I_{xx}$ versus $I_{xy}$ at a three-terminal junction. A negative voltage is applied to the left contact, while the other contacts are grounded. $I_{xx}$ is devided according to Kirchhoff’s current rule. Neglecting the partial reflection of current at ohmic contacts, only one egde is negatively charged, other edges are at ground potential. $I_{xy}$ is guided by the charged edge and flows entirely from the left to the lower contact.
In our experiments, we measure the current, which corresponds to the integral of the current densities in Eq. (II.2) across the Hall bar. For the partition of the current at a junction we can now use the following rules: The ohmic $I_{xx}$ branches according to Kirchhoffs current law. For simplicity, we neglect the small resistance $R_{0}$ of the 2DES ($R_{i}/R_{0}\sim 50$). The non-ohmic $I_{xy}$ partitions according to Eq. (4), if we include non-integer filling factors [$R_{\text{H}}=(\nu G_{0})^{-1}$]. Therefore, in the limit $\mu B\gg 1$, the prediction for the current division of the Drude model becomes identical to that of our coherent and chiral model derived in Sec. II.1. To illustrate the different behaviors of $I_{xx}$ and $I_{xy}$, in Fig. 3 we sketch the current division at a three-terminal junction. Combining both contributions, the Drude model predicts
$I_{i}$ $=$ $I_{xx}^{(i)}+I_{xy}^{(i)}\quad\text{with}$
$I_{xx}^{(i)}$ $=$ $\frac{1}{1+(\mu B)^{2}}\,\frac{1}{R_{i}}\left[\frac{\sum_{m=1}^{n}V_{m}/R_{m}}{\sum_{m=1}^{n}1/R_{m}}-V_{i}\right]$ (10)
$I_{xy}^{(i)}$ $=$ $\frac{(\mu B)^{2}}{1+(\mu B)^{2}}\,\frac{1}{R_{i}}\left[\frac{1}{P_{i}}\left(\frac{S_{n}}{P_{n}-1}+S_{i}\right)-V_{i}\right]\,$
for the current flowing into the $i$th contact.
@@
The original scope of the classical Drude model was a description of diffusive transport at $\mu B\lesssim 1$. Interestingly, in the limit $\mu B\to\infty$, the first term of Eq. (II.2) vanishes, $I_{xx}\to 0$, and the prediction of the Drude model becomes identical to Eq. (4). This property also means that for the limits $\mu B\lesssim 1$ versus $\mu B\gg 1$, Eq. (II.2) precisely describes both of our limiting models. For intermediate $\mu B$, Eq. (II.2) smoothly interpolates between the limits. This is useful, because $\mu(B)$ strongly oscillates due to the Shubnikov-de-Haas oscillations.
@@
Multiterminal current measurements yield the distribution of current into various contacts. In this section, we discuss the results of multi-terminal measurements and compare them with theory predictions in a wide magnetic field region corresponding to $1\lesssim\nu<\infty$. We will see, that even though multiterminal measurements do not provide the local current distribution across the Hall bar, their comparison with reasonable limiting models can provide answers to the questions about chirality of the current for the QHE.
@@ III.1 Sample and measurement setup
Figure 4: Sketch of a Hall bar with six contacts, characterized by their ohmic resistances, $R_{i}$. We apply a voltage $V_{1}$ to contact 1 and measure currents into contacts 1, 3, and 5. Contacts 2, 4, and 6 can be left floating or connected to electrical ground. Current amplifiers at contacts 3 and 5 result in additional voltages of $\simeq 20\mu$V applied to contacts 3 and 5 (not shown, but taken into account for generating model curves). $\vec{B}$ points into the plane of the 2DES.
It is etched from a GaAs / (Al,Ga)As heterostructure, which contains a two-dimensional electron system (2DES) 130 nm beneath the surface. We performed the experiments at a temperature of $T\simeq 300\,$mK in a He-3 evaporation cryostat. At this temperature and zero magnetic field, the electron density and mobility of the 2DES, determined from Hall measurements, are $n_{\textrm{s}}\simeq 1.2\times 10^{11}\textrm{ cm}^{-2}$ and $\mu\simeq 3.95\times 10^{6}\textrm{cm}^{2}/\textrm{Vs}$ (yielding a mean-free-path of $\lambda_{\text{m}}\simeq 23\,\mu$m), respectively.

### v3

@@ II.1 Coherent and chiral model of n-terminal device
Figure 2: Simplified circuit diagram of an $n$-terminal device. The inner circle indicates coherent and chiral current flow inside the 2DES. Each contact k=1,,nk=1,...,n has an ohmic resistance $R_{k}$ and a voltage $V_{k}$ applied in respect to the electrical ground. $V_{k}^{\prime}$ are the voltages resulting at the intersections between the ohmic contacts and the coherent inner circle. The magnetic field vector pointing into the plane of the 2DES gives rise to a right handed chirality of a free charge which translates into a left handed chirality of the device, cf. circled arrow.
The inner circle in Fig. 2 indicates the 2DES, the radial lines symbolize the $n$ terminals k=1,2,,nk=1,2,...,n connected via ohmic contacts to the 2DES. External voltages $V_{k}$ can be applied to the contacts in respect to ground, while $V_{k}^{\prime}=V_{k}+I_{k}R_{k}$ are the voltages present at the intersects of the 2DES and each contact. The current $I_{k}$ flows through the $k$th contact via the ohmic resistance $R_{k}$ to ground.
@@
where $G_{kj}$ are the conductance coefficients between mutual contacts. They can be formally derived from the scattering matrix of the problem. The left-hand side of Eq. (2) is a general formulation of the transmission properties of the 2DES based on the Landauer-Büttiker formalism, while the last term of Eq. (2) applies Ohm’s law to the macroscopic contacts. If the transport within the 2DES is chiral, only those $G_{k,j\neq k}$ connecting adjacent contacts can be non-zero. The direction of the chirality is linked to the direction of the magnetic field vector. In our case sketched in Fig. 2, we consider $\vec{B}$ pointing into the plane of the 2DES, such that $G_{kj}\neq 0$ for $k=j+1$. Assuming coherent transport for the inner circle of Fig. 2, the scattering matrix is unitary, implying $\sum_{k}G_{kj}=\sum_{j}G_{kj}=0$, such that our coherent and chiral conductivity tensor takes the form
[…]
@@
where $G_{0}=e^{2}/h$ and $\nu$ is the filling factor of the (spin-resolved) Landau-levels. The assumption of coherence, which guarantees unitarity of the scattering matrix and the validity of Eq. (3) is expected for the absence of scattering inside an ICS. For the quantized plateaus of the Hall resistance, $\nu$ is an integer. The Eq. (3) might be considered as our formal definition of chiral current flow. Inserting Eq. (3) into Eq. (2) yields a solvable linear equation system providing us with the prediction of the current into the $k$th contact
$I_{k}=\frac{1}{R_{k}}\left[\frac{1}{P_{k}}\left(\frac{S_{n}}{P_{n}-1}+S_{k}\right)-V_{k}\right]\,$, (4)
where we used $V_{k}^{\prime}=\frac{1}{P_{k}}\left(\frac{S_{n}}{P_{n}-1}+S_{k}\right)$ with
$P_{k}=\left\{\begin{matrix}\prod_{j=1}^{k}\left(1+\frac{1}{\nu G_{0}R_{j}}\right)\,;&k=1,2,...,n\\[4.30554pt] 1\,;&k=0\phantom{,2,...,n}\end{matrix}\right$. (5)
$\text{and }S_{k}=\sum_{j=1}^{k}P_{j-1}\frac{V_{j}}{\nu G_{0}R_{j}}\,.$ (6)
@@
In Appendix A we provide a derivation of Eq. (4). A simplifying special case of some relevance is that of $V_{1}\neq 0$ but all other applied voltages $V_{i>1}=0$, which yields the current ratios
$\frac{I_{k>1}}{I_{1}}=\frac{{P_{n}}/{P_{k}}}{1-{P_{n}}/{P_{1}}}\,\frac{1}{\nu G_{0}R_{k}}\,$ (7)
irrespective of the sign of $V_{1}$. To visualize the nature of chiral current flow, in Fig. 3
Figure 3: Chiral current flow in a y-junction: Simplifying sketches for a three-terminal junction for $\vec{B}$ pointing into the plane. In panel (a) a negative, in (b) a positive voltage $V_{1}$ is applied to contact 1, while the other contacts are grounded. Contact resistances are neglected, such that charge accumulates at the edge between contacts 1 and 2, only. Due to the direction of the applied magnetic field current flows clockwise (electrons counter clockwise). Bright blue arrows indicate the path way of the current generated by $V_{1}$. For $V_{1}<0$ it flows from contact 2 to 1; for $V_{1}>0$ it flows from 1 to 3 and then to 2. The faint arrows indicate a persistent chiral current that flows still for $V_{1}=V_{2}=V_{3}=0$.
we sketch the chiral current path predicted by our model for just three terminals with $V_{2}=V_{3}=0$, while we neglect the contact resistances for simplicity. Arrows indicate the technical current direction, the electrons flow in the opposite direction. The clockwise chiral current flow is fixed by the direction of $\vec{B}$ pointing from the surface downwards through the 2DES plane. Contacts 2 and 3 are grounded, while $V_{1}<0$ in panel (a) and $V_{1}>0$ in (b). The chiral current predicted by Eq. (3) includes a persistent contribution carried by states below the chemical potentials of all three leads, which flows even for $V_{1}=V_{2}=V_{3}=0$. Such a persistent current is expected for the plateaus of the QHE [26, 35]. The faint arrows indicate persistent current, while the bright blue arrows indicate the flow of the sum of the persistent current and the additional current imposed by applying $V_{1}\neq 0$. For $V_{1}<0$, excess electrons move directly from contact 1 to contact 2 ($1\to 2$) according to $V_{1}<V_{2}=V_{3}=0$, hence, the imposed current flows from $2\to 1$. For $V_{1}>0$, the number of electrons flowing from $2\to 3$ and that from $3\to 1$ are equal, while the number of electrons flowing from $1\to 2$ is smaller. Therefore, the imposed current flows from $1\to 3\to 2$.
In summary, the path of the chiral current is fixed by the direction of the magnetic field. As a result, if we neglect contact resistances, all current generated by applying a voltage to contact 1, always flows through contact 2 independent of the sign of $V_{1}$ or the number of additional grounded contacts. Upon adding contact resistances, Eq. (4) predicts corrections, such as some current flowing through additional contacts.
@@ II.2 Drude model of an n-terminal device
where the specific conductivity (defined at $B=0$) is $\sigma_{0}=en_{\text{s}}\mu$. For a 2DES laterally confined in the $y$-direction, in the steady state we expect $j_{y}=0$, which implies $E_{y}=-\mu BE_{x}$ and, consequently, $j_{x}=\sigma_{0}E_{x}=R_{\text{H}}^{-1}E_{y}$.
@@
The longitudinal component $j_{xx}=\sigma_{xx}E_{x}$ is the ohmic term, proportional to the specific conductivity and driven by the longitudinal electric field $E_{x}$. Caused by the additional Hall resistance, at finite $B$ the ohmic current density is reduced by the factor $\left[1+(\mu B)^{2}\right]^{-1}$ and vanishes for $\mu B\gg 1$. The transversal component $j_{xy}=\sigma_{xy}E_{y}$ is proportional to the perpendicular electric field. For $\mu B\gg 1$, with $j_{xy}\to E_{y}/R_{\text{H}}$ it becomes independent of the ohmic resistance. This implies, that for $\mu B\gg 1$ momentum scattering is absent. At the same time, the transversal $j_{xy}$ follows the charged edge of the Hall bar. This way, it is bent into a current carrying contact.
In our experiments, we measure the current, which corresponds to the integral across the Hall bar of the current densities in Eq. (II.2). To illustrate the contributions of $I_{xx}$ and $I_{xy}$, in Fig. 4
@@
Figure 4: Diffusive current flow in a y-junction (Drude model): Simplifying sketches showing the partitioning of the current contributions $I_{xx}$ versus $I_{xy}$ in a three-terminal junction for $\vec{B}$ pointing into the plane. In panel (a) a negative, in (b) a positive voltage is applied to contact 1, while the other contacts are grounded. Contact resistances are neglected, such that charge accumulates at the edge between contacts 1 and 2, only. $I_{xx}$ is devided according to Kirchhoff’s current law. $I_{xy}$ is guided by the charged edge and flows perpendicular to the Hall field $E_{y}$ (indicated by dashed arrows). $I_{xy}$ flows from contact 2 to 1 (opposite to electron flow).
we sketch the current division according to the Drude model in a three-terminal junction for the identical conditions as in Fig. 3, where we sketched current path ways within the chiral model, ($V_{1}\neq 0$, $V_{2}=V_{3}=0$, $R_{1}=R_{2}=R_{3}=0$). $I_{xx}$ branches according to Kirchhoff’s current law, while the non-ohmic $I_{xy}$ is bent by the edge charges (respective magnetic field) into the next available (non-floating) contact in current direction. The classical Drude model predicts unidirectional current flow, which is in no way chiral. Nevertheless, the current component $I_{xy}$ is bent such that it flows between contacts 1 and 2, as does the imposed current predicted by our chiral model in Sec. II.1.
Applying the Drude model to our general $n$-terminal sample allows us to predict the currents $I_{k}$ flowing into each contact in the diffusive regime. If we assume a homogeneous 2DES in our Hall bar and neglect possible asymmetries in the contact configuration, we may incorporate the resistance $R_{0}(B)$ of the 2DES into the $n$ contact resistances, and find
$I_{k}$ $=$ $I_{xx}^{(k)}+I_{xy}^{(k)}\quad\text{with}$
$I_{xx}^{(k)}$ $=$ $\frac{1}{1+(\mu B)^{2}}\,\frac{1}{R_{k}}\left[\frac{\sum_{j=1}^{n}V_{j}/R_{j}}{\sum_{j=1}^{n}1/R_{j}}-V_{k}\right]$ (10)
$I_{xy}^{(k)}$ $=$ $\frac{(\mu B)^{2}}{1+(\mu B)^{2}}\,\frac{1}{R_{k}}\left[\frac{1}{P_{k}}\left(\frac{S_{n}}{P_{n}-1}+S_{k}\right)-V_{k}\right]\,$
for the current flowing into the $k$th contact. In Appendix Sec. B we provide a detailed explanation of this result and discuss the relevance of $R_{0}(B)$ for the case of an asymmetric configuration of contacts.
@@
The original scope of the classical Drude model was a description of diffusive transport at $\mu B\lesssim 1$. However, in the limit $\mu B\to\infty$, the first term of Eq. (II.2) vanishes, $I_{xx}\to 0$, and the prediction of the Drude model becomes identical to Eq. (4), see Appendix B for a discussion. Therefore, we can interpret Eq. (II.2) as follows: for the limit $\mu B\lesssim 1$, which occurs for non-integer $\nu$ between the plateaus of the QHE, Eq. (II.2) describes diffusive transport based on the Drude model. For the opposite limit $\mu B\gg 1$, which occurs for integer $\nu$ on the plateaus, Eq. (II.2) precisely describes transport according to our chiral model. For intermediate $\mu B$, Eq. (II.2) smoothly interpolates between the limits. This is useful, because $\mu(B)$ strongly oscillates due to the SdH oscillations.
@@
Multiterminal current measurements yield the distribution of current into various contacts. In this section, we discuss the results of such experiments and compare them with theory predictions in a wide magnetic field range corresponding to filling factors between $1\lesssim\nu<\infty$.
@@ III.1 Sample and measurement setup
[…]
For multi-terminal current measurements, we apply the voltage of $V_{1}=\pm 1\,$mV to the source contact 1 (using a Keithley 2450 source meter). We measure the current $I_{1}(B)$ flowing from the Hall bar into contact 1, using the source meter, and the currents $I_{3}(B)$ and $I_{5}(B)$ flowing into contacts 3 or 5, respectively, using Basel Precision Instr. current amplifiers. Contacts 3 and 5 are connected to the electrical ground, while the remaining contacts 2, 4, and 6 may be either electrically floating or connected to ground depending on the experiment. For a first impression, in Fig. 6
Figure 6: Currents $I_{k}(B)$ measurement of a symmetric y-junction: Contacts 2, 4 and 6 in Fig. 5 are floating, $V_{1}=+1\,$mV for $B>0$ and $V_{1}=-1\,$mV for $B<0$, contacts 3 and 5 are connected to the electrical ground. Measured currents are $I_{1}$ (black line), $I_{3}$ (red) and $I_{5}$ (blue), gray lines correspond to $I_{1}$. We find $I_{3}>I_{5}$ for $B>0$ and $I_{5}>I_{3}$ for $B<0$, where $I_{1}+I_{3}+I_{5}=0$.
we present the measured currents $I_{1}(B)$, $I_{3}(B)$ and $I_{5}(B)$ for two three-terminal measurements (contacts 2, 4 and 6 are floating, such that contacts 3 and 5 correspond to contacts 2 and 3 in the sketches in Fig. 3 and Fig. 4). We define $B>0$ for $\vec{B}$ pointing downwards (from the surface through the 2DES) and $B<0$ for $\vec{B}$ pointing upwards. We applied the voltages $V_{1}=+1\,$mV and swept $B>0$ versus $V_{1}=-1\,$mV for $B<0$. In both cases, for $B=0$, the currents $I_{3}\simeq I_{5}$, because the contact resistances are almost equal. (Of course, the sign of the currents depends on the sign of the voltage.) As expected, $I_{1}=I_{3}+I_{5}$ always $|I_{1}|$ decreases as $|B|$ is increased because of $R_{\text{H}}\propto|B|$. Importantly, for $B>0$, we observe $|I_{3}|>|I_{5}|$, while for $B<0$, we find $|I_{5}|>|I_{3}|$. This behavior is independent of the sign of $V_{1}$ (not shown here). These data directly visualize our prediction that the current distribution in the various grounded contacts is governed by the direction of $\vec{B}$, only, cf. Eq. (II.2) or the sketches in Fig. 3 and Fig. 4.
[…]
contact $k$ resistance $R_{k}$ ($\Omega$) voltage $V_{k}$ (mV) geometry $a_{k}$
1 2420 1.000 1
2 2420 0.5
3 2400 -0.021 1
4 2350 1
5 2420 -0.018 1
6 2400 0.5
Table 1: Parameters of the circuit sketched in Fig. 5. Contact resistances $R_{k}$ include ohmic contacts, cables and RC-filters. Voltage $V_{1}=1$ mV is applied to contact 1, $V_{3}$ and $V_{5}$ are offset voltages of the current amplifiers connected to the respective contacts. Contacts 2, 4 and 6 are either floating or directly connected to ground. Hence, $V_{2}$, $V_{4}$ and $V_{6}$ are zero (variable), if the respective contact is grounded (floated). To account for the contribution of the Hall bar itself to the resistance, we defined a geometry parameter and use $R_{k}\to R_{k}+a_{k}/[n_{\text{s}}e\mu(B)]$ in the model calculations of the ohmic contributions; $a_{2}$ and $a_{6}$ are smaller, as they are closest to contact 1, such accounting for a shorter current path that comes into effect when the respective contacts are grounded. The correction is negligible for $B=0$.
This preparation allows us to perform quantitative comparisons between measurements and models. For the reminder of the article we discuss experiments with $V_{1}=1\,$mV and $B>0$. For the chiral model, electrons then move from contact to contact along the counterclockwise pathway $1\to 2\to 3\to 4\to 5\to 6\to 1$. Since the external voltage that we apply to contact 1 exceeds all other voltages by far (and all contact resistances are similar), for a sizable magnetic field $B>0$, according to the prediction of both models and our first experiment shown in Fig. 6, the largest current will flow between contacts 1 and the next grounded contact following in counterclockwise direction.
III.2 Three-terminal measurements
We start our quantitative comparison between measurements and models for the three-terminal measurements already presented in Fig. 6 above. Here, the grounded contacts 3 and 5 are arranged in a mirror symmetric geometry in respect to the horizontal axis through contact 1. In Fig. 7(a),